# Safe Deep Policy Adaptation

### v2

@@ VI EXPERIMENTS
In this section, we demonstrate extensive experimental results in Inverted Pendulum, Safety Gym [41], and Real-World RC Car, to address the following questions:
@@ VI-A Inverted Pendulum
We adopt an inverted pendulum simulation augmented with external wind perturbations from [42]. Wind (with fixed magnitude, and variable angle from 0 to 360 degrees) is introduced as the variable environment configuration $e_{t}$. In this setting, we aim to evaluate the adaptive safety of SafeDPA, and its dual capability in safety and task performance.
@@ VI-B Safety Gym Simulation Experiments
To answer Q2, we implement SafeDPA on the Safety Gym [41] benchmark. The point robot in Safety Gym has higher state dimensions $x_{t}\in\mathbb{R}^{17}$, compared to pendulum’s $x_{t}\in\mathbb{R}^{2}$. We introduce an external force $F_{t}\in\mathbb{R}^{2}$ (on the 2D plane) within the MuJoCo [43], and define $F_{t}$ as environment configuration. The control task is to navigate the point robot to a target and avoid hazards. For this setting, we also incorporate standard RL methods (PPO [34] and TRPO [35]) and end-to-end safe RL methods including CPO [19], PPO-Lagrangian, and TRPO-Lagrangian [18] as benchmarks. Similar to experiments in Section VI-A, we evaluate the algorithms across various environments with the direction of $F_{t}$ ranging from 00 to $360$ degrees. The results are demonstrated in Figure 5. Remarkably, our proposed SafeDPA is the sole algorithm that maintains both high performance and safety across all $F_{t}$ values. In contrast, other baseline algorithms tend to excel in either performance or safety, but not in both aspects simultaneously.

### v3

@@ VI EXPERIMENTS
In this section, we demonstrate extensive experimental results in Inverted Pendulum, Safety Gym [Ray2019], and Real-World RC Car, to address the following questions:
@@ VI-A Inverted Pendulum
We adopt an inverted pendulum simulation augmented with external wind perturbations from [shi2021meta]. Wind (with fixed magnitude, and variable angle from 0 to 360 degrees) is introduced as the variable environment configuration $e_{t}$. In this setting, we aim to evaluate the adaptive safety of SafeDPA, and its dual capability in safety and task performance.
@@ VI-B Safety Gym Simulation Experiments
To answer Q2, we implement SafeDPA on the Safety Gym [Ray2019] benchmark. The point robot in Safety Gym has higher state dimensions $x_{t}\in\mathbb{R}^{17}$, compared to pendulum’s $x_{t}\in\mathbb{R}^{2}$. We introduce an external force $F_{t}\in\mathbb{R}^{2}$ (on the 2D plane) within the MuJoCo [todorov2012mujoco], and define $F_{t}$ as environment configuration. The control task is to navigate the point robot to a target and avoid hazards. For this setting, we also incorporate standard RL methods (PPO [schulman2017proximal] and TRPO [schulman2015trust]) and end-to-end safe RL methods including CPO [achiam2017constrained], PPO-Lagrangian, and TRPO-Lagrangian [tessler2018reward] as benchmarks. Similar to experiments in Section VI-A, we evaluate the algorithms across various environments with the direction of $F_{t}$ ranging from 00 to $360$ degrees. The results are demonstrated in Figure 5. Remarkably, our proposed SafeDPA is the sole algorithm that maintains both high performance and safety across all $F_{t}$ values. In contrast, other baseline algorithms tend to excel in either performance or safety, but not in both aspects simultaneously.